A sample contains 4.80 grams of O2 gas has an initial volume of 15.0 L. If pressure and temperature remain constant: What is the final volume if 0.500 mole of O2 gas is added?
1
Expert's answer
2020-03-30T10:50:27-0400

We calculate the molar volume of gas under given conditions:


"n(O_2) = m\/M = 4.80\/32 = 0.15mol"

"V_m = V\/n = 15.0\/0.15 = 100 L\/mol"

Then, the amount of oxygen under given conditions:


"V(O_2) = V_mn = 100*0.5 = 50.0L"

Total:


"V(O_2)_{total} = 50.0 +15.0 = 65.0L"

Answer: 65.0L